if
2+3=10
7+2=63
6+5=66
8+4=96
WHAT DOES 9+7=

Answer by Earlsdon(6294)

If:
2+3 = 10
7+2 = 63
6+5 = 66
8+4 = 96 Then...
9+7 = 144